OpenStudy (anonymous):

Let f(x) = 5x + 12. Find f-1(x).

OpenStudy (anonymous):

@johnweldon1993 WHAT I know this is different but WHAT

OpenStudy (johnweldon1993):

lol yeah this one is COMPLETELY different...it's an inverse! similar looking right? f(-1) = plug in -1 for 'x' and f-1(x) = inverse equation So what you do... y = 5x + 12 You switch the 'x' and the 'y' x = 5y + 12 And now you use algebra to solve for 'y' again

OpenStudy (anonymous):

Yep you lost me which one do I use

OpenStudy (johnweldon1993):

lol f-1(x) here...the inverse... so your equation is y = 5x + 12 (Note how I replaced f(x) with y ...same thing) What you do to solve an inverse equation *f-1(x)* is you switch the position of the 'x' and the 'y' in the equation... x = 5y + 12 Now you need to solve for 'y' again... x = 5y + 12 Lets subtract 12 from both sides x - 12 = 5y And then divide everything by 5 \[\large y = \frac{x - 12}{5}\] and that would be your answer...does that make sense?

OpenStudy (anonymous):

After looking at it a few times I think so... ahah. There's a couple others I will try on my own right now.

OpenStudy (johnweldon1993):

If you need help feel free to message me :)